# Towards Continual Expansion of Data Coverage: Automatic Text-guided Edge-case Synthesis

### v1

@@ Appendix D Analysis of bias in data splits
Appendix E Custom evaluation metric proposal
E.1 Background and motivation
@@
E.2 Limitations of existing analysis methods
@@ E.3 Proposal: mAP w/o TP
The core idea is to first define non-weak instances as the set of annotations that the baseline Model $\mathcal{A}$ already detects correctly (its true positives (TP)). Then, when evaluating Model $\mathcal{B}$, we exclude its correct predictions on these non-weak instances from the calculation. This focuses the evaluation squarely on the subset of data that was challenging for the baseline model, making any improvements on these blind spots much more apparent.
E.4 Step-by-Step calculation
@@
    • $\mathcal{G}_{\text{weak}}$: All other annotations, which correspond to Model $\mathcal{A}$’s FN or were not associated with any high-confidence prediction.
@@
We provide an example of the filtered ground-truth annotations and predictions in Figure 4.
E.5 Significance and benefits
@@
Appendix F Analysis in a class-wise manner
@@
Appendix G Analysis in a tag-wise manner
@@
Appendix H Analysis on transferability
To assess whether a dataset deemed effective for one model remains valid for another, a core concern in data-centric evaluation, we conducted transferability experiments. This approach aligns with recent findings [12, 7] that emphasize robust evaluation designs, acknowledging that instance informativeness may vary across model configurations and necessitating experimental methods to identify broadly transferable data. To examine the transferability of synthetic data across model scales, we trained YOLOv11-medium and YOLOv11-large using datasets generated by the proposed pipeline with YOLOv11-small. As summarized in Table 4, 5, the automatic still achieves the highest mAP and mAP w/o TP.
Table 4: YOLOv11-medium performance with different training dataset configurations.
Train dataset Count mAP mAP w/o TP
train-D 3,187 0.366 $\pm$ 0.004 -
train-D + naive v0 19,122 0.426 $\pm$ 0.002 -
train-D + naive v0 + naive v1 35,057 0.429 $\pm$ 0.004 0.415 $\pm$ 0.005
train-D + naive v0 + manual v1 35,057 0.428 $\pm$ 0.005 0.414 $\pm$ 0.005
train-D + naive v0 + automatic v1 35,057 0.429 $\pm$ 0.003 0.416 $\pm$ 0.005
Table 5: YOLOv11-large performance with different training dataset configurations.
Train dataset Count mAP mAP w/o TP
train-D 3,187 0.371 $\pm$ 0.001 -
train-D + naive v0 19,122 0.419 $\pm$ 0.001 -
train-D + naive v0 + naive v1 35,057 0.431 $\pm$ 0.003 0.414 $\pm$ 0.004
train-D + naive v0 + manual v1 35,057 0.429 $\pm$ 0.007 0.411 $\pm$ 0.008
train-D + naive v0 + automatic v1 35,057 0.432 $\pm$ 0.005 0.415 $\pm$ 0.006
Appendix I Analysis of rephrasing patterns
(a) naive
(b) manual
(c) automatic
Figure 5: Comparison of naive, manual, and automatic by UMAP [21] of CLIP [28] embeddings. Gray points represent real data, while the colored ones indicate synthetic data, with colors closer to yellow denoting higher density.
@@ Appendix I Analysis of rephrasing patterns
Appendix J Examples of synthetic data
Each row in Figure 6 corresponds to the same real image serving as the base caption, with synthetic examples generated using different approaches: naive, manual, and automatic. This row-wise arrangement allows for a direct comparison of how each generation strategy affects the resulting pseudo annotations. While the automatic does not necessarily produce more instances than the manual variant, it exhibits a higher diversity in backgrounds, capturing variations in lighting conditions, camera angles, and scene context. In contrast, manual, although increasing the number of difficult cases, often produces more repetitive scenes with constrained background variation.
real naive manual automatic
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Comparison of real and synthetic data. Pseudo annotations are marked with different colors indicating different classes (Bus, Bike, Car, Pedestrian, Truck). Best viewed in color.
Appendix K Analysis of edge-ness
Table 6: Comparison of mean and median statistics of the edge-ness values
Dataset Mean Median
train-D 2.211 $\pm$ 0.056 2.315 $\pm$ 0.173
naive v0 2.485 $\pm$ 0.208 2.587 $\pm$ 0.032
naive v1 2.624 $\pm$ 0.002 2.555 $\pm$ 0.005
manual v1 2.709 $\pm$ 0.004 2.649 $\pm$ 0.005
automatic v1 2.724 $\pm$ 0.011 2.656 $\pm$ 0.007
We empirically validate the core mechanism of our pipeline: its ability to generate images with high edge-ness. We measure the training loss of the discriminative model on data generated by each method. As shown in Table 6, data from our proposed pipeline consistently induces the highest mean and median loss. This confirms that our preference learning framework successfully steers the LLM to generate captions that translate into verifiably more difficult examples for the discriminative model, thereby creating effective training signals for improving robustness.
Appendix L Prompt templates
@@
L.1 Image captioning prompt
@@
L.2 Rephrasing prompt by the manual baseline
@@
L.2.1 Rephrasing prompt to construct preference dataset
@@
L.2.2 Rephrasing prompt with the preference-tuned LLM

### v2

@@ Appendix C Analysis of bias in data splits
Appendix D Custom evaluation metric proposal
D.1 Background and motivation
@@
D.2 Limitations of existing analysis methods
@@ D.3 Proposal: mAP w/o TP
The core idea is to first define non-weak instances as the set of annotations that the baseline Model $\mathcal{A}$ already detects correctly (its true positives (TP)). Then, when evaluating Model $\mathcal{B}$, we exclude its correct predictions on these non-weak instances from the calculation. This focuses the evaluation squarely on the subset of data that was challenging for the baseline model, making any improvements on these blind spots much more apparent. We provide an example of the filtered ground-truth annotations and predictions in Figure 7.
D.4 Step-by-Step calculation
@@
    • $\mathcal{G}_{\text{weak}}$: Other annotations, which correspond to Model $\mathcal{A}$’s FN or were not associated with any high-confidence prediction.
@@
D.5 Significance and benefits
@@
Appendix E Analysis in a class-wise manner
@@
Appendix F Analysis in a tag-wise manner
@@
Appendix G Analysis of rephrasing patterns
@@
Appendix H Prompt templates
@@
H.1 Image captioning prompt
@@
H.2 Rephrasing prompt by the manual baseline
@@
H.2.1 Rephrasing prompt to construct preference dataset
@@
H.2.2 Rephrasing prompt with the preference-tuned LLM